OpenStudy (finishwithan):

Help please:) I have an idea but I don't know if it is right..I will give medal:)

OpenStudy (finishwithan):

A box contains 100 colored chips; some are pink and some are blue. Milton chooses a chip at random, records the color, and places it back in the bag. Milton has recorded 12 pink chips and 18 blue chips. Using these results, what is the predicted number of pink chips in the box? 30 40 82 88

OpenStudy (finishwithan):

i think it is either 30 or 40..but i do not exactly know:(

OpenStudy (griefer101):

12+18=30

OpenStudy (pigsruleyeah):

30

OpenStudy (finishwithan):

i was thinking that but it is starting to not make sense since he put them back into the bag after he counted them...

OpenStudy (loser66):

nope

OpenStudy (loser66):

12 pink 18 blue, that is \(\frac{12pink}{18blue}=\dfrac{2pink}{3blue}\)

OpenStudy (loser66):

So, on 100 chips, you divide it into 5 parts, each part is 20 chips. the pink occupies 2 parts, that is ???

OpenStudy (loser66):

what??? 2*20 =10?

OpenStudy (finishwithan):

i meant 40..i did 20/2 oops

OpenStudy (loser66):

ok

OpenStudy (finishwithan):

wait @Loser66 ...so it is 40?

OpenStudy (loser66):

yup

OpenStudy (finishwithan):

thanks!